OpenStudy (anonymous):

Solve: 1.) 2 log 2 x - log 2 (x+3) = 2 2.)Solve: log 4 x = 1/2 3.)log b 9 = 2 4.)log 2 y = 1/3 (log 2 27)

OpenStudy (anonymous):

you type that right?

OpenStudy (anonymous):

the 2 is sopposed to be to the 2nd power

OpenStudy (anonymous):

hmmm And the log thing?

OpenStudy (anonymous):

I am good with the Solve for X... then a crap tone of BS so if you need just brake it down for me to help you

OpenStudy (anonymous):

I'm assuming some of these numbers are called the base of the log, which is the smaller number that goes right after log. These are all based on the properties of logs, which you have probably learned. 1. Logs with the same base: their innards can be multiplied together. 2(x)(x+3)=2 Solve. 2. Base to the answer equals innards. 4^(1/2) = x Anything to the 1/2 is always just the square root, so the square root of 4 is 2. 3. Base to the answer equals innards. b^2 = 9 Square root both sides, b=3. 4. Logs with same base divided by each other: numerator innards - denominator innards. Divide both sides by log (base 2) 27. log (base 2) y / log (base 2) 27 = 1/3 Subtract innards. y - 27 = 1/3 Solve.